OpenStudy (rosales0228):

Tina want to save money for school. Tina invest 1,000 in an account that pays and interest of 8%. NEED HELP PLEASE!!!! GIVING FAN AND MEDAL!!! How many years will it take for the account to reach 1,300 round your answer to the nearest hundredth.

OpenStudy (anonymous):

If Tina receives simple interest, it will take her 3.75 years. $1,300-$1,000 = $300 $300/($1,000*.08) = 3.75 years If Tina receives compound interest, it will take her 3.41 years. \[$1,300= $1,000(1+.08)^{n}\]\[1.3=1.08^{n}\]\[\log(1.3)=\log(1.08)^{n} \]\[.1139=n*.0334\]\[n=3.41\]